OpenStudy (anonymous):

please solve....find the equation of the line satisfying the conditions given. Express the answer in standard form. Through (2,-2) parallel to -9x +8y=82

OpenStudy (anonymous):

y=(9x/8)-4.25

OpenStudy (anonymous):

Slope of parallel line: m = 9/8 y -2 = (9/8) ( x+2) 8( y -2) = 9 (x +2) => 9x - 8y = -34